A resultant force at 200N is applied to a vehicle of mass 400kg which is initially at rest on a horizontal surface at the total for opposing the motion is constant at 800N.1.Calculate the acceleration of the vehicle, 2.Kinetic energy of the vehicle at 5second After the force is applied? 3.Total power develop 5second after the force first applied?
1
Expert's answer
2020-02-24T11:03:13-0500

1.


"a=\frac{R}{m}=\frac{200}{400}=0.5\frac{m}{s^2}"

2.


"K=0.5mv^2=0.5m(at)^2=\\=0.5(400)(0.5\cdot 5)^2=1250\ J"

3.

"P=0.5Fv=0.5(R+F_f)at"

"P=0.5(800+200)(0.5)(5)=1250\ W"